OpenStudy (anonymous):

A savings account compounds interest, at a rate of 15%, once a year. Elizabeth puts $800 in the account as the principal. How can Elizabeth set up a function to track the amount of money she has? @idku

OpenStudy (ranga):

She starts with $800. After 1 year, the interest will be 800 * .15. Add the principal of $800 and you have 800 + 800*.15 or 800*1.15 in her account. After two years the account will have 800*1.15*1.15 = 800*1.15^2 SO after t years the account will have: 800*1.15^t

OpenStudy (ranga):

\[\Large A(t) = 800(1.15)^t\] Where A is the amount after t years.

OpenStudy (anonymous):

A(x) = 800(1 + .15)x where .15 is the interest rate @ranga